OpenStudy (anonymous):

compute the derivative y=log10 (5x^4-2/x^3)

OpenStudy (paxpolaris):

\[\Large y= \log_{10}\left( 5x^4-2\over x^3 \right)\] this can be simplified to:\[\Large y= \log_{10}\left( 5x^4-2 \right)-\log_{10} \left(x^3\right)\]

OpenStudy (paxpolaris):

find the dervative using chain rule , and using:\[\large \log_{10}(b)={\ln(b)\over \ln(10)}\]

OpenStudy (anonymous):

no Its 2 over X^3 not he whole thing .\[y=\log_{10} \left( 5x ^{4} -2\div x^3\right)\]

OpenStudy (paxpolaris):

if: \(\Large y=\log_{10}\left( f \left( x \right) \right)\) using chain rule: \[\large y'={1 \over \ln(10)}\times \left( 1\over f(x) \right)\times f'(x)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you ! May God bless you !